Jump to main content.


Drake Chemical

EPA Daily Situation Reports for August 1998

U. S. ARMY CORPS OF ENGINEERS
DRAKE CHEMICAL SUPERFUND SITE

182 MYRTLE STREET
LOCK HAVEN, PA 17745
PHONE 717-748-4505
FAX 717-748-7549

Monday, August 31, 1998

FOR DATES: From 7:00 a.m. Sunday, Aug. 30, 1998 to 7:00 a.m. Monday, Aug. 31, 1998. WEATHER: Sunny and clear. The temperature range was 62oF to 82oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 962.3 tons of contaminated soil was thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.2 tons per hour while on line. Down time was contributed to daily weigh belt calibration. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, stockpiling, feed preparation, lime mixing, debris handling, debris separation, rock cleaning, rock backfilling, ash handling, and ash backfilling. Continued real time perimeter air monitoring.

The contractor backfilled the remainder of the clean burn material above elevation 545. The material was a mixture of clay and small rock that was imported from off site. The material had been temporarily stockpiled across the railroad tracks east of the site.

Received bottom ash analytical report (bot-ash-08-26-98-202) for bin 3. Test results indicated that the thermally treated soil met the project requirement for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None noted.

PREPARED BY: Chad Thompson for David Modricker (USACE)


Sunday, August 30, 1998

FOR DATES: From 7:00 a.m. Saturday, Aug. 29, 1998 to 7:00 a.m. Sunday, Aug. 30, 1998.

WEATHER: Sunny and clear. The temperature range was 62oF to 78oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.6 hours and 949.6 tons of contaminated soil was thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.2 tons per hour while on line. Down time was contributed to weigh belt calibration and a flame out of the SCC burners. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, stockpiling, feed preparation, lime mixing, debris handling, debris separation, ash handling, and ash backfilling. Continued real time perimeter air monitoring and collected the weekly OSHA 93 and NIOSH 5518 samples.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handing, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Richard Lupold for David Modricker (USACE)


Saturday, August 29, 1998

FOR DATES: From 7:00 a.m. Friday, Aug. 28, 1998 to 7:00 a.m. Saturday, Aug. 29, 1998.

WEATHER: Sunny and clear. The temperature range was 59oF to 85oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 962.5 tons of contaminated soil was thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.3 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, stockpiling, feed preparation, lime mixing, debris handling, debris separation, ash handling, and ash backfilling. Continued real time perimeter air monitoring (PAM) and began the weekly modified OSHA method 93 testing for BNA at the PAM and offsite air monitoring stations. Received bottom ash analytical report (bot-ash-08-24-98-200) for bin 1. Test results indicated that the thermally treated soil met project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handing, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Sandy Downs (PADEP)

PREPARED BY: Steve Gibson for David Modricker (USACE)


Friday, August 28, 1998

FOR DATES: From 7:00 a.m. Thursday, Aug. 27, 1998 to 7:00 a.m. Friday, Aug. 28, 1998.

WEATHER: Sunny and clear. The temperature range was 58oF to 85oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 962.7 tons of contaminated soil was thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.3 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, mass excavation, stockpiling, feed preparation, lime mixing, debris handling, debris separation, rock cleaning, rock backfilling, ash handling, and ash backfilling. Continued real time perimeter air monitoring.

Received bottom ash analytical report (bot-ash-08-22-98-198) for bin 6. Test results indicated that the thermally treated soil met the project requirement for unrestricted backfilling on-site. Received bottom ash analytical report (bot-ash-08-23-98-199) for bin 7. Test results indicated that the thermally treated soil met project requirements for restricted backfilling (i.e. above elevation 553) due to the concentration of chromium being greater than the PADEP drinking water standard, but less than 25 times the standard. Received fly ash analytical report (fly-ash-08-22-98-036) for bin 'D'. Test results indicated that the fly ash materials met project requirements for restricted backfilling (i.e. above elevation 553) due to the concentrations of chromium and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handing, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Sandy Downs (PADEP)

PREPARED BY: Steve Gibson for David Modricker (USACE)


Thursday, August 27, 1998

FOR DATES: From 7:00 a.m. Wednesday, Aug. 26, 1998 to 7:00 a.m. Thursday, Aug. 27, 1998.

WEATHER: Overcast, clearing in the afternoon. The temperature range was 55oF to 85oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 959.0 tons of contaminated soil was thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.1 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, mass excavation, stockpiling, feed preparation, lime mixing, debris handling, debris separation, ash handling, and ash backfilling. Continued real time perimeter air monitoring. Received bottom ash analytical report (bot-ash-08-21-98-197A) for bin 5. Test results indicated that the thermally treated soil met project requirements for restricted backfilling (i.e. above elevation 553) due to the concentration of arsenic being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handing, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for David Modricker (USACE)


Wednesday, August 26, 1998

FOR DATES: From 7:00 a.m. Tuesday, Aug. 25, 1998 to 7:00 a.m. Wednesday, Aug. 26, 1998.

WEATHER: Humid with an evening thunderstorm. The temperature range was 64oF to 76oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.0 hours and 917.1 tons of contaminated soil were thermally treated during the time frame of this report. Off-line time is attributed primarily to manual feed stoppage as a precaution during local thunderstorm activity. The TDF processed contaminated soil at an average rate of 39.9 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, stockpiling, feed preparation, lime mixing, debris handling, debris separation, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities and began a sampling event for Total Suspended Particles (TSP) and metals at the perimeter air monitoring (PAM) stations.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handing, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Sandy Downs, Steve Webster (PADEP)

PREPARED BY: Steve Gibson for David Modricker (USACE)


Tuesday, August 25, 1998

FOR DATES: From 7:00 a.m. Sunday, Aug. 23, 1998 to 7:00 a.m. Monday, Aug. 24, 1998.

WEATHER: Humid with an early afternoon thunderstorm. The temperature range was 63oF to 84oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 22.6 hours and 895.2 tons of contaminated soil was thermally treated during the time frame of this report. Off-line time is attributed primarily to manual feed stoppage as a precaution during local thunderstorm activity. The TDF processed contaminated soil at an average rate of 39.6 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, stockpiling, feed preparation, lime mixing, debris handling, rock cleaning, rock backfilling, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities. Two rolloffs of PPE (202 and 240) were shipped off site for disposal at a hazardous waste landfill. Received bottom ash analytical reports (bot-ash-08-17-98-194, bot-ash-08-18-98-195, bot-ash-08-20-98-196) for bottom ash bins 2, 3, and 4. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547). Received fly ash analytical report (fly-ash-08-16-98-035) for fly ash bin C. Analytical test results indicated that the fly ash materials met the project requirements for restricted backfilling (i.e. above elevation 553) due to the concentrations of cadmium, chromium, and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handing, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: David Polish, Hollis Sloggin, Richard Kuhn, Kimberly Lane (USEPA)

PREPARED BY: Steve Gibson for David Modricker (USACE)


Monday, August 24, 1998

FOR DATES: From 7:00 a.m. Sunday, Aug. 23, 1998 to 7:00 a.m. Monday, Aug. 24, 1998.

WEATHER: Morning fog, partly cloudy. The temperature range was 64oF to 84oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility(TDF). The TDF was on-line for 23.8 hours and 956.6 tons of contaminated soil was thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 40.2 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, stockpiling, feed preparation, lime mixing, ash handling, and ash backfilling. Recovered OSHA 93 and NIOSH 5518 samples. Continued real time perimeter air monitoring activities.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handing, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None.

PREPARED BY: Danny Zion for David Modricker (USACE)


Sunday, August 23, 1998

FOR DATES: From 7:00 a.m. Saturday, Aug. 22, 1998 to 7:00 a.m. Sunday, Aug. 23, 1998.

WEATHER: Partly cloudy and humid with intermittent showers in the afternoon. The temperature range was 67oF to 85oF.

I. CURRENT SITUATION: The contractor is operating the Thermal Destruction Facility (TDF) 24 hours a day, 7 days a week. The TDF was on-line for 23 hours, 57 minutes and 958.7 tons of contaminated soil was thermally treated during the time frame of this report. Off-line time was attributed to weigh belt calibration. The TDF processed contaminated soil at an average rate of 40.0 tons per hour while on line. Related ongoing site activities include feed preparation.

II. ACTIONS TAKEN: The contractor continued TDF operations, stockpiling, feed preparation, ash handling, and ash backfilling. Started 24 hour time weighted average NIOSH 5518 and OSHA 93 sampling activities. Continued real time perimeter air monitoring activities.

III. FUTURE PLANS: Future plans include operation of the TDF, continuation of excavation, drum handing, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Richard Lupold for David Modricker (USACE)


Saturday, August 22, 1998

FOR DATES: From 7:00 a.m. Friday, Aug. 21, 1998 to 7:00 a.m. Saturday, Aug. 22, 1998.

WEATHER: Overcast and clearing in the Afternoon. The temperature range was 57oF to 83oF.

I. CURRENT SITUATION: The contractor is operating the Thermal Destruction Facility (TDF) 24 hours a day, 7 days a week. The TDF was on-line for 21.5 hours and 858.6 tons of contaminated soil was thermally treated during the time frame of this report. Off-line time was attributed to additional maintenance being performed on the connections between the newly installed feed screws and the feed screw motor. The TDF processed contaminated soil at an average rate of 39.9 tons per hour while on line. Related ongoing site activities include feed preparation.

II. ACTIONS TAKEN: The contractor continued TDF operations, stockpiling, feed preparation, debris separation, lime mixing, washing large concrete debris for backfilling, rock cleaning, rock backfilling, ash handling, and ash backfilling. The contractor cleaned out the sediment in sediment basin number 2. Continued real time perimeter air monitoring activities.

III. FUTURE PLANS: Future plans include operation of the TDF, continuation of excavation, drum handing, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for David Modricker (USACE)


Friday, August 21, 1998

Report not available.


Thursday, August 20, 1998

FOR DATES: From 7:00 a.m. Wednesday, Aug. 18, 1998 to 7:00 a.m. Thursday, Aug. 19, 1998.

WEATHER: Clear and mild. The temperature range was 47oF to 73oF.

I. CURRENT SITUATION: Waste feed is currently off to the Thermal Destruction Facility (TDF). Waste feed was stopped at 10:30 a.m. on Wednesday, August 19, 1998. The TDF is on hot hold while the contractor prepares to replace worn feed screws and performs maintenance on the TDF. Waste feed is expected to resume Friday, August 21 or Saturday, August 22, 1998. The TDF was on-line for 3.5 hours and 104.7 tons of contaminated soil was thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 29.9 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations and began preventative maintenance on the ID fan, replacement of the feed screw drive motor and drive chain, and the replacement of the feed screws. During the TDF down time, the contractor performed the monthly calibrations of the TDF monitoring equipment and conducted monthly interlock checks. At 3:27 p.m., the Thermal Relief Valve (TRV) was opened as part of the interlock checks. Prior to the opening of the TRV, feed to the incinerator had been shut down for approximately 5 hours and the contractor notified the appropriate authorities.

The contractor continued WWTP operations, excavation, stockpiling, feed preparation, debris separation, lime mixing, rock cleaning, rock backfilling, ash handling, and ash backfilling. The contractor placed odor suppressant foam on the excavation face. Continued real time perimeter air monitoring activities.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handing, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: The contractor could not install the modified feed screws due to some minor deformation of the shaft. Pending machine work on the shaft scheduled for today, the modified feed screws will be installed.

V. NOTED VISITORS: Fred Castillio (OHM corporate engineer), Dave Mummert (OHM CIH)

PREPARED BY: Steve Gibson for David Modricker (USACE)


Wednesday, August 19, 1998

FOR DATES: From 7:00 a.m. Tuesday, Aug. 18, 1998 to 7:00 a.m. Wednesday, Aug. 19, 1998.

WEATHER: Overcast with periods of rain. The temperature range was 50oF to 77oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line 23.4 hours and 751.6 tons of contaminated soil were thermally treated during the time frame of this report. Off-line time was attributed to multiple stoppages of the feed screws. The TDF processed contaminated soil at an average rate of 32.1 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, WWTP operations, excavation, stockpiling, feed preparation, debris separation, lime mixing, rock cleaning, rock backfilling, ash handling, and ash backfilling. The contractor placed odor suppressant foam on the excavation face. Continued real time perimeter air monitoring activities.

III. FUTURE PLANS: A maintenance shutdown for feed screw replacement is scheduled to start about 11 a.m. today (August 19, 1998). A planned TRV opening is scheduled to occur about 3 p.m. today. Major activities for the next week include operation of the TDF, continuation of excavation, drum handing, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Sandy Downs, Larry Newcomer, Steve Webster (PADEP)

PREPARED BY: Steve Gibson for David Modricker (USACE)


Tuesday, August 18, 1998

FOR DATES: From 7:00 a.m. Monday, Aug. 17, 1998 to 7:00 a.m. Tuesday, Aug. 18, 1998.

WEATHER: Overcast with light rain. The temperature range was 68oF to 75oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line 22.9 hours and 807 tons of contaminated soil was thermally treated during the time frame of this report. Off-line time was attributed to multiple stoppages of the feed screws. The TDF processed contaminated soil at an average rate of 35.2 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, WWTP operations, excavation, feed preparation, debris separation, lime mixing, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities. Received bottom ash analytical report (bot-ash-08-11-98-189) for bin 4. The chemical analysis indicated that the thermally treated soil met the project requirements for restricted backfilling on-site due to the concentration of selenium being greater than the PADEP drinking water standard, but less than 25 times the standard. Received fly ash analytical report (fly-ash-08-12-98-034) for bin B. The chemical analysis indicated that the fly ash materials met the project requirements for restricted backfilling on-site due to the concentrations of chromium and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard. Both materials are restricted to backfill above elevation 553.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handing, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE)


Monday, August 17, 1998

FOR DATES: From 7:00 a.m. Sunday, Aug. 16, 1998 to 7:00 a.m. Monday, Aug. 16, 1998.

WEATHER: Cloudy, rain Monday morning. The temperature range was 67oF to 83oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line 21.68 hours and 832.8 tons of contaminated soil was thermally treated during the time frame of this report. Off-line time was primarily for repair of the feed screws drive chain. The TDF processed contaminated soils at an average rate of 38.4 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities. Received bottom ash analytical report (bot-ash-08-11-98-188) for bin 3. The chemical analysis indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handing, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None

PREPARED BY: Danny Zion for David Modricker (USACE)


Sunday, August 16, 1998

FOR DATES: From 7:00 a.m. Saturday, Aug. 15, 1998 to 7:00 a.m. Sunday, Aug. 16, 1998.

WEATHER: Overcast with evening rain. The temperature range was 66oF to 80oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line 20.58 hours and 794 tons of contaminated soil was thermally treated during the time frame of this report. Off-line time was attributed to numerous short term feed screw stoppages, maintenance on the feed screws drive chain, repair of transformer by PP&L, and a SCC flame out. The TDF processed contaminated soils at an average rate of 38.6 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handing, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None

PREPARED BY: Richard Lupold for David Modricker (USACE)


Saturday, August 15, 1998

FOR DATES: From 7:00 a.m. Friday, Aug. 14, 1998 to 7:00 a.m. Saturday, Aug. 15, 1998.

WEATHER: Overcast with evening rain. The temperature range was 65oF to 78oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line 23.4 hours and 919 tons of contaminated soil were thermally treated during the time frame of this report. Off-line time was attributed to several short term feed screw stoppages. The TDF processed contaminated soils at an average rate of 39.3 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, WWTP operations, mass excavation, feed preparation, lime mixing, debris separation, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities. In accordance with ASTDR regulations, the contractor's QC department deployed Summa canisters as a result of a low level NMOC reading recorded by PAM station 3. The contractor further investigated the situation and was not able to detect any action level readings from the site's PAM stations, no further action was taken. Received bottom ash analytical report (bot-ash-08-10-98-187) for bin 2. The chemical analysis indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handing, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Sandy Downs (PADEP)

PREPARED BY: Steve Gibson for David Modricker (USACE)


Friday, August 14, 1998

FOR DATES: From 7:00 a.m. Thursday, Aug. 13, 1998 to 7:00 a.m. Friday, Aug. 14, 1998.

WEATHER: Clear. The temperature range was 55oF to 81oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line 23.5 hours and 931.1 tons of contaminated soil were thermally treated during the time frame of this report. Off-line time was attributed to stopped feed screws. The TDF processed contaminated soils at an average rate of 39.6 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, WWTP operations, feed preparation, lime mixing, debris separation, ash handling, and ash backfilling. Excavation in the exclusion zone was restarted on this date. Continued real time perimeter air monitoring activities. Received bottom ash analytical report (bot-ash-08-09-98-186) for bin 1. The chemical analysis indicated that the thermally treated soil met the project requirements for restricted backfilling on-site (i.e. above elevation 553) due to the level of cadmium being greater than the PADEP drinking water standard, but less than 25 times the standard. Received fly ash analytical report (fly-ash-08-07-98-033) for bin A. The chemical analysis indicated that the fly ash materials met the project specifications for restricted backfilling on-site (i.e. above elevation 553) due to the levels of selenium and chromium being greater than the PADEP drinking water standared, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handing, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Dr. S. Tucker (OHM)

PREPARED BY: Steve Gibson for David Modricker (USACE)


Thursday, August 13, 1998

FOR DATES: From 7:00 a.m. Wednesday, Aug. 12, 1998 to 7:00 a.m. Thursday, Aug. 13, 1998.

WEATHER: Overcast, clearing later. The temperature range was 55oF to 81oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line 20 hours and 786 tons of contaminated soil were thermally treated during the time frame of this report. The off-line time is attributed to the stoppage of feed in order to replace the thermocouples in the kiln. The TDF processed contaminated soils at an average rate of 39.3 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, WWTP operations, feed preparation, lime mixing, debris separation, ash handling, and ash backfilling. Began grading a staging area within the exclusion zone in preparation for excavation of the south side of the stockpile next week. Much of the material excavated was observed to be demolition debris. Continued real time perimeter air monitoring activities. Received bottom ash analytical report (bot-ash-08-08-98-185) for bin 7. The chemical analysis indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handing, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Dr. S. Tucker (OHM), Sandy Downs (PADEP).

PREPARED BY: Steve Gibson for David Modricker (USACE)


Wednesday, August 12, 1998

FOR DATES: From 7:00 a.m. Tuesday, Aug. 11, 1998 to 7:00 a.m. Wednesday, Aug. 12, 1998.

WEATHER: Hazy and hot. The temperature range was 63oF to 86oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line 23.6 hours and 873.2 tons of contaminated soil was thermally treated during the time frame of this report. Off-line time is attributed to feed screw stoppages. OHM is investigating the cause of the feed screw stoppages, which do influence the feed rate. The TDF processed contaminated soils at an average rate of 37.0 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, WWTP operations, feed preparation, lime mixing, debris separation, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities. Received bottom ash analytical report (bot-ash-08-07-98-183) for bin 6. The chemical analysis indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handing, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Dr. S. Tucker (OHM), Sandy Downs, Steve Webster, Larry Newcomer, Mike Welch (PADEP).

PREPARED BY: Steve Gibson for David Modricker (USACE)


Tuesday, August 11, 1998

FOR DATES: From 7:00 a.m. Monday, Aug. 10, 1998 to 7:00 a.m. Tuesday, Aug. 11, 1998.

WEATHER: Overcast with periods of heavy rain. The temperature range was 65oF to 76oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line 22.9 hours and 900.7 tons of contaminated soil was thermally treated during the time frame of this report. Off-line time is attributed to feed screw stoppages. The TDF processed contaminated soils at an average rate of 39.3 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, WWTP operations, feed preparation, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities, began the thermally controlled modified OSHA 93 24-hour air sampling for BNA at the PAM and off-site air monitoring stations. Began the 24-hour BNA sampling at the off-site air monitoring stations by use of the NIOSH 5518 method. Received bottom ash analytical reports (bot-ash-08-05-98-182 and bot-ash-08-06-98-183) for bins 4 and 5. The chemical analysis indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handing, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for David Modricker (USACE)


Monday, August 10, 1998

FOR DATES: From 7:00 a.m. Sunday, Aug. 9, 1998 to 7:00 a.m. Monday, Aug. 10, 1998.

WEATHER: Partly cloudy and warm. The temperature range was 65oF to 82oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 20 hours, 56 min. and 847.3 tons of contaminated soil was thermally treated during the time frame of this report. The TDF was off-line for 2 hours and 58 minutes due to an electrical storm. The TDF processed contaminated soils at an average rate of 40.5 tons per hour. Related ongoing site activities include feed preparation.

II. ACTIONS TAKEN: Continued TDF operations, feed preparation, dust control, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities. Received ash analytical report for bot-ash-08-4-98-181. The chemical analysis indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, continuation of excavation, drum handing, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: NONE

V. NOTED VISITORS: NONE

PREPARED BY: Danny Zion for David Modricker (USACE)


Sunday, August 9, 1998

FOR DATES: From 7:00 a.m. Saturday, Aug. 8, 1998 to 7:00 a.m. Sunday, Aug. 9, 1998.

WEATHER: Mostly sunny and warm. The temperature range was 60oF to 82oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23 hours, 40 min. and 958.9 tons of contaminated soil were thermally treated during the time frame of this report. The TDF processed contaminated soils at an average rate of 40.5 tons per hour. Related ongoing site activities include feed preparation.

II. ACTIONS TAKEN: Continued TDF operations, feed preparation, dust control, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handing, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: NONE

V. NOTED VISITORS: NONE

PREPARED BY: Rick Lupold for David Modricker (USACE)


Saturday, August 8, 1998

FOR DATES: From 7:00 a.m. Friday, Aug. 7, 1998 to 7:00 a.m. Saturday, Aug. 8, 1998.

WEATHER: Clear and hot. The temperature range was 58oF to 87oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.93 hours and 969.0 tons of contaminated soil were thermally treated during the time frame of this report. The TDF processed contaminated soils at an average rate of 40.5 tons per hour. Related ongoing site activities include feed preparation.

II. ACTIONS TAKEN: Continued TDF operations, feed preparation, debris separation, dust control, rock cleaning, rock backfilling, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handing, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Sandy Downs (PADEP).

PREPARED BY: Steve Gibson for David Modricker (USACE)


Friday, August 7, 1998

FOR DATES: From 7:00 a.m. Thursday, Aug. 6, 1998 to 7:00 a.m. Friday, Aug. 7, 1998.

WEATHER: Clear and hot. The temperature range was 55oF to 87oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 965.8 tons of contaminated soil were thermally treated during the time frame of this report. The TDF processed contaminated soils at an average rate of 40.6 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, WWTP operations, feed preparation, debris separation, dust control, rock cleaning, rock backfilling, ash handling, and ash backfilling. The contractor shipped two rolloffs of PPE and miscellaneous non-hazardous debris to a hazardous waste landfill. Continued real time perimeter air monitoring activities. Deployed Summa air collection canisters at PAM stations 1 and 3.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handing, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Sandy Downs (PADEP)

PREPARED BY: Steve Gibson for David Modricker (USACE)


Thursday, August 6, 1998

FOR DATES: From 7:00 a.m. Wednesday, Aug. 5, 1998 to 7:00 a.m. Thursday, Aug. 6, 1998.

WEATHER: Clear and hot. The temperature range was 63oF to 84oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 970.1 tons of contaminated soil were thermally treated during the time frame of this report. The TDF processed contaminated soils at an average rate of 40.6 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, WWTP operations, feed preparation, debris separation, dust control, rock cleaning, rock backfilling, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handing, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for David Modricker (USACE)


Wednesday, August 5, 1998

FOR DATES: From 7:00 a.m. Tuesday, August 4, 1998 to 7:00 a.m. Wednesday, August 5, 1998.

WEATHER: Clear. The temperature range was 61oF to 86oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 967.4 tons of contaminated soil were thermally treated during the time frame of this report. The TDF processed contaminated soils at an average rate of 40.5 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, WWTP operations, feed preparation, debris separation, dust control, rock cleaning, rock backfilling, ash handling, and ash backfilling. Continued real time perimeter air monitoring and began the 24-hour modified OSHA-93 sampling event for BNA at the on-site and off-site air monitoring stations.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handing, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Sandy Downs, Steve Webster, Mike Welch, Larry Newcomer (PADEP).

PREPARED BY: Steve Gibson for David Modricker (USACE)


Tuesday, August 4, 1998

FOR DATES: From 7:00 a.m. Monday, August 3, 1998 to 7:00 a.m. Tuesday, August 4, 1998.

WEATHER: Partly cloudy. The temperature range was 56oF to 84oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 22.9 hours and 912.2 tons of contaminated soil were thermally treated during the time frame of this report. Off-line time was attributed to repairing the weigh belt speed sensor. While on line, the TDF processed contaminated soils at an average rate of 39.8 tons per hour. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF operations, WWTP operations, feed preparation, debris separation, dust control, rock cleaning, rock backfilling, ash handling, and ash backfilling. Continued real time perimeter air monitoring and off-site air monitoring activities.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handing, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Sandy Downs (PADEP)

PREPARED BY: Steve Gibson for David Modricker (USACE)


Monday, August 3, 1998

FOR DATES: From 7:00 a.m. Sunday, August 2, 1998 to 7:00 a.m. Monday, August 3, 1998.

WEATHER: Clear. The temperature range was 53oF to 83oF.

I. CURRENT SITUATION: The contractor is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). After the week long shut down for maintenance, contaminated soil feed was initiated at 0952 hours. However, after feed was started, a series of start up difficulties occurred with the unit that led to the unit being off-line for a part of the day/night. Difficulties causing the down time have been attributed to a damaged weigh belt speed sensor, which has been repaired. The TDF was on-line for 16.3 hours and an estimated 565 tons of contaminated soil were thermally treated during the time frame of this report.

II. ACTIONS TAKEN: Continued TDF operations, feed preparation, dust control, ash handling, and ash backfilling. Continued real time perimeter air monitoring activities. Started monthly perimeter air sampling event for volatiles and metals.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF; continuation of excavation, drum handling, drum bulking, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination

IV. PROBLEMS NOTED OR ANTICIPATED: None except as described in item I above.

V. NOTED VISITORS: None.

PREPARED BY: Danny Zion for David Modricker (USACE)


Sunday, August 2, 1998

FOR DATES: From 7:00 a.m. Saturday, August 1, 1998 to 7:00 a.m. Sunday, August 2, 1998.

WEATHER: Sunny. The temperature range was 50oF to 79oF.

I. CURRENT SITUATION: The Thermal Destruction Facility (TDF) is currently heating up in order to resume the thermal treatment of contaminated soils. Feed is tentatively scheduled to begin Sunday morning. The contractor continues to work 24 hours a day, operating and maintaining the TDF. The contractor is performing related activities such as soil feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued feed preparation, feed stockpiling, lime mixing, debris separating, dust control, rock and inorganic debris cleaning, rock backfilling at the bottom of the excavation, and backfilling of clean ash. Continued decontamination of equipment scheduled to be removed from the site. Continued real time perimeter air monitoring activities.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning, rock backfilling, and inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Danny Zion for Dave Modricker (USACE)


Saturday, August 1, 1998

FOR DATES: From 7:00 a.m. Friday, July 31, 1998 to 7:00 a.m. Saturday, August 1, 1998.

WEATHER: Partly cloudy to clear. The temperature range was 49oF to 78oF.

I. CURRENT SITUATION: The Thermal Destruction Facility (TDF) is currently heating up in order to resume the thermal treatment of contaminated soils. Feed is tentatively scheduled to begin Sunday morning. The contractor continues to work 24 hours a day, operating and maintaining the TDF. The contractor is performing related activities such as soil feed preparation and debris separation.

II. ACTIONS TAKEN: Continued TDF maintenance, which included the installation of the new feed screws, installation of a new feed screw cooling jacket, and the installation of the reconditioned Induced Draft (ID) fan motor. In addition to the TDF maintenance, the TDF instruments and gauges were calibrated prior to start up. The rotary kiln burners were lit at approximately 1705 hours.

The contractor continued feed preparation, feed stockpiling, lime mixing, debris separating, dust control, rock and inorganic debris cleaning, and rock backfilling at the bottom of the excavation. The contractor removed the storm inlet at the south end of the site. Continued real time perimeter air monitoring activities.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, rock cleaning, rock backfilling, and inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for Dave Modricker (USACE)


Region 3 | Mid-Atlantic Cleanup | Mid-Atlantic Superfund |EPA Home | EPA Superfund Homepage


Local Navigation


Jump to main content.